NEW LINE LISTS INCLUDING INTENSITIES FOR THE C_{2} SWAN SYSTEM (d^{3}Π_{g}-a^{3}Π_{u}), C_{2} SINGLET SYSTEMS AND ROVIBRATIONAL TRANSITIONS WITHIN THE NH X^{3}Σ^{−} GROUND STATE.

\underline{J. S. A. BROOKE}, Department of Chemistry, University of York, York, UK; P. F. BERNATH, Department of Chemistry & Biochemistry, Old Dominion University, Norfolk, VA, USA; C. M. WESTERN, School of Chemistry, University of Bristol, Bristol, UK; T. W. SCHMIDT and G. B. BACSKAY, School of Chemistry, The University of Sydney, New South Wales, Australia; M. C. VAN HEMERT, Department of Chemistry, Gorlaeus Laboratories, Leiden University, The Netherlands; G. C. GROENENBOOM, Theoretical Chemistry, Institute for Molecules and Materials (IMM), Radboud University Nijmegen, Nijmegen, The Netherlands.

Line lists including both positions and absolute intensities are required by those who wish to calculate abundances from spectroscopic observations. New line lists for the diatomic molecules C_{2} and NH are presented. Line strengths for the C_{2} Swan system (d^{3}Π_{g}-a^{3}Π_{u}) have been calculated for vibrational bands with v′=0-10 and v′′=0-9, and J values up to J =34-96, based on previous observations in 36 vibrational bands. Line positions from several sources were combined with the results from recent deperturbation studies of the v′=4 and v′=6 levels, to provide updated molecular constants. The line strengths are based on a recent ab initio calculation of the transition dipole moment function. Einstein A coefficients and f -values were also calculated for the vibrational bands of the Swan system. A line list has been made available, including observed and calculated line positions, Einstein A coefficients and oscillator strengths (f -values). This list will be useful for astronomers, combustion scientists and materials scientists who utilize C_{2} Swan spectra. Similar work is being carried out for a number of singlet systems of C_{2}, in which a single global fit of positions will be performed and intensities calculated. For NH, a line list is being created for the infrared transitions within the X^{3}Σ^{−} ground state, including previous observations of rotational lines within v=0-2, and rovibrational lines of the Δv=1 sequence up to v′=6. The intensity calculations use a recently calculated dipole moment function which has been used to calculate the lifetime of the v=1 level, but was unpublished. A line list like that for the C_{2} Swan system will be made available, which will be useful for astronomers investigating NH and nitrogen abundance in cool objects.